OpenStudy (anonymous):

Find the three geometric means between 2 and 162.

OpenStudy (anonymous):

6,18,54

OpenStudy (welshfella):

yes - those are the correct answers

OpenStudy (welshfella):

the numbers 2,6,18,54,162 ar a geometric series 2 is the first term and 162 is the 5th ( as there are 3 geometric means ) therefore r^2 = 162/2 = 81 r is the 4th root of 81 so the series is 2 , 2*3 , 2*3^2, 2*3^3 , 2* 3^4 = 2, 6 , 18, 54 , 162 6 , 18 and 54 are the 3 geometric means

OpenStudy (welshfella):

* that is r^4 = 162/2 = 81 ( on the 4th line) the 4th root of 81 is 3

OpenStudy (anonymous):

oh i kinda see now